OpenStudy (anonymous):

Explain how raising a quotient to a power is similar to raising a product to a power.

OpenStudy (anonymous):

maybe this: \[(\frac{ a }{ b })^2 = \frac{ a^2 }{ b^2 }\] and \[(ab)^2 = a^2 b^2\]

OpenStudy (anonymous):

i should have used a variable letter instead of 2, but holds for any number